Two RL circuits are connected by the mutual inductance L12. Find the currents I1 and I2 flowing in the left and the right circuits if the left circuit is driven by a time dependent emf V Voeiut. Set both currents as I1(t) 21e (wt-01) and (t) 12e iut-on) and find the amplitude ratio i1 and the phase difference D1 p2 (20pt). 12 11 22
Answer #1

impedence of circuit 1

Z = R +i \omega L11

current I1(t) = V/z = Vo exp(i\omega t) / ( R + i \omega L11 )

                                 = Vo/ ( R2 +\omega2 L211 ) exp(i\omega t) *( R - i \omega L11 )

                                 = Vo/ ( R2 +\omega2 L211 ) exp(i\omega t) *exp(- i\phi1   ) , where tan\phi1 = \omega L11 /R ,

                                 = i1 exp i( \omega t -\phi1 ) , i1 = Vo/ ( R2 +\omega2 L211 )

mutual inductance = L12

emf in circuit 2

V2(t) =L12 d I1(t) /dt = L12 i \omega i1exp i( \omega t -\phi1 )

current in circuit 2:

I2(t) = V2/ (R+i L22) = L12 i \omega i1exp i( \omega t -\phi1 ) (R - i L22) /(R2+ \omegaL222)

                               = L12\omega i1exp i( \omega t -\phi1 ) ( L22 - i R ) /(R2+ L222)

                              = L12\omega i1exp i( \omega t -\phi1 ) exp(- i\phi2 )/(R2+ \omega2L222) , where tan\phi2 = R/\omegaL22

                             = i2 exp i( \omega t -\phi1 - \phi2 ) where i2 = L12\omega i1/(R2+ \omega2L222)

i1/i2 = (R2+ \omega2L222)/L12    

phase difference

(\phi1 - \phi2) = arcTan(\omega L11 /R ) - arcTan(R/\omegaL22 )